A charge, q1, of +17 micro-coulombs is located at x = 0, y = 5.3 cm, a charge q2 of -42 micro-coulombs is located at x = 0, y = -4.0 cm, an unknown charge of q3, is located at x =- 4.9 cm, y = 0 cm, and a charge, q4, -8 micro-coulombs is located at x = +18.2 cm, y = 0 cm. The direction of the total electric force on q4 is 135.0 degrees from the +x axis, counter-clockwise. What is the unknown charge, q3, in micro-coulombs? If negative, include a negative sign in your answer.
Answer #1

1 17x 10-6 C r14 F1 r1 0.053 m F2 135 40.182 m - r0.049 m r2 0.04 m r24 2-42x 106C

Using Pythagorean theorem

r14 = sqrt(r12 + r42) = sqrt((0.053)2 + (0.182)2) = 0.1896 m

r24 = sqrt(r22 + r42) = sqrt((0.04)2 + (0.182)2) = 0.1863 m

also

\theta1 = tan-1(r1/r4) = tan-1(0.053/0.182) = 16.24 deg

\theta2 = tan-1(r2/r4) = tan-1(0.04/0.182) = 12.4 deg

r34 = r3 + r4 = 0.049 + 0.182 = 0.231 m

F1 = force by charge q1 on q4 = k q1 q4/r142 = (9 x 109) (17 x 10-6) (8 x 10-6)/(0.1896)2 = 34.05 N

Fi = - (F1 Cos\theta1) i + (F1 Sin\theta1) j = - (34.05 Cos16.24) i + (34.05 Sin16.24) j = - 32.7 i + 9.5 j

F2 = force by charge q2 on q4 = k q2 q4/r242 = (9 x 109) (42 x 10-6) (8 x 10-6)/(0.1863)2 = 87.13 N

F2 = (F2 Cos\theta2) i + (F2 Sin\theta2) j = (87.13 Cos12.4) i + (87.13 Sin12.4) j = 85.1 i + 18.7 j

F = magnitude of net force

\underset{F}{\rightarrow}= F Cos135 i + F Sin135 j = - (0.71) F i + (0.71) F j

F3 = magnitude of force by charge q3 on q4

\underset{F_{3}}{\rightarrow} = F3 i

net force is given as

\underset{F}{\rightarrow}= Fi + F2 + \underset{F_{3}}{\rightarrow}

- (0.71) F i + (0.71) F j = (- 32.7 i + 9.5 j ) + ( 85.1 i + 18.7 j) + (F3 i)

comparing the coefficients of "j" on both side

(0.71) F = 9.5 + 18.7

F = 39.1 N

comparing the coefficients of "i" on both side

- 0.71 F = - 32.7 + 85.1 + F3

- 0.71 (39.1) = - 32.7 + 85.1 + F3

F3 = - 80.2 N

since the force F3 is negative , hence force by charge q3 comes out to be towards negative x-axis. hence charge q3 is positive

F3 = k q3 q4/r342

80.2 = (9 x 109) q3 (8 x 10-6)/(0.231)2

q3 = 59.4 x 10-6 C